# GranQ: Granular Zero-Shot Quantization with Channel-Wise Activation Scaling in QAT

### v4

@@ 5.3.2 Efficiency Evaluation
Table 5: Ablation study of vectorization in scaling computation and quantization. GranQ applies vectorization to both stages, achieving high accuracy with low latency. Latency is measured with batch size 16. Experiments use AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
@@
While fine-grained activation quantization often incurs high computational cost due to per-channel scaling, GranQ maintains efficiency by vectorizing the scaling computation. By reshaping activation tensors, our method eliminates the overhead of iterative channel-wise operations. To assess this in practice, we measured quantization latency across various batch sizes (16, 32, 64, 128, and 200). Figure 5 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels (Algorithm 1 in appendix).
@@
The two channel-wise methods perform per-channel quantization. The conventional method computes scaling factors sequentially, whereas GranQ vectorizes the scaling computation across channels for improved efficiency. A key observation from Figure 5 is that GranQ incurs only a slight increase in latency compared to conventional layer-wise quantization. With only a minor increase in latency, GranQ achieves higher accuracy than the layer-wise baseline, as shown in Table 5. Compared to scalar-loop-based channel quantization, which implements the same computation sequentially, GranQ achieves nearly identical accuracy. Furthermore, Table 2 shows that GranQ substantially reduces quantization error and more effectively preserves activation information. These results demonstrate that GranQ enables fine-grained quantization with significantly improved efficiency. Additional analysis and experimental results on inference efficiency are provided in Table 7 of the appendix.
6 Discussion
6.1 Why GranQ is Effective?
@@
The most common approaches to activation quantization (i.e., layer-wise quantization) are known for their low computational cost and fast processing speed. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 2). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. These limitations become more severe in low-bit settings, especially when synthetic data exhibit distributional shifts.
To overcome these issues, GranQ computes and stores per-channel scaling factors for each input in a unified vector form, enabling parallel processing across all channels. Unlike conventional methods, it performs quantization based on each input’s channel-wise distribution, resulting in more precise and efficient computation.
6.2 Limitation and Future-work
In ZSQ settings, synthetic data often exhibit broader and more distorted activation distributions compared to original data (Figure 6 in appendix). GranQ addresses this challenge by applying dynamic quantization for each channel based on its activation distribution, making it particularly effective in QAT scenarios. However, GranQ is less suitable for PTQ, which relies on static activation quantization. As future work, we plan to develop quantization methods that preserve the fine-grained scaling of GranQ while enhancing adaptability to PTQ settings.
7 Conclusion
Layer-wise and channel-wise quantization are both commonly used in weight quantization, as weights remain fixed after training. In contrast, channel-wise scaling in activation quantization is more challenging, as activation values vary with each input, making it difficult to apply fine-grained scaling. These limitations become more pronounced in QAT settings, where large amounts of data are used. Under these conditions, quantization is strongly affected by activation distributions. This issue is especially critical in ZSQ, where training data are synthesized using internal statistics from the FP model. Synthetic data generated in this way often produce distorted activation distributions compared to original data, resulting in increased quantization error during scaling. To address these issues, we propose GranQ, a dynamic channel-wise quantization method that adapts scaling factors to each input’s activation distribution. This enables more precise quantization even under distorted distributions, as often observed in ZSQ. In addition, GranQ improves computational efficiency by applying per-input, per-channel scaling through vectorized operations, which is rarely adopted in activation quantization due to its implementation complexity.
Through extensive experiments, we demonstrate that GranQ achieves superior performance while preserving computational efficiency. Compared with various SOTA methods, GranQ demonstrates competitive performance and achieves notable improvements. Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
@@ Appendix A Appendix
A.1 Notation
We provide an overview of the main symbols and definitions employed throughout the paper. It specifically includes essential elements such as the original activation ($A$), quantized activation ($A_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
@@ A.1 Notation
$A$ Original activation value
$A_{q}$ Quantized activation value
$A_{\text{deq}}$ Dequantized activation value
$A_{\min}$, $A_{\max}$ Minimum and maximum activation values
$Q$ Quantization operator
$s$ Scaling factor
$z$ Zero-point for quantization
@@
$\min\limits_{h,w}(A)$, $\max\limits_{h,w}(A)$ Channel-wise min/max activations
$\mathbb{R}^{C\times H\times W}$ Tensor of shape $C\times H\times W$ in $\mathbb{R}$
[…]
$\|\cdot\|_{2}$ L2 norm (Euclidean norm)
$\in$ Set membership ("element of")
@@
Forward Quantization $x_{q}=\left\lfloor\frac{x}{s}+z\right\rceil$
Dequantization $x_{\text{deq}}=s(x_{q}-z)$
Scaling Factor $s=\frac{x_{\max}-x_{\min}}{2^{b}-1}$
Zero-point $z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$
@@
Table 6: Notation for activation quantization equations, algorithm, and evaluation metrics.
A.2 Algorithm
While prior works apply vectorized computation during the quantization step, they typically compute per-channel scaling parameters using iterative channel-wise operations. GranQ differs by also vectorizing the scaling computation, allowing both steps to be processed in parallel. Algorithm 1 outlines the complete process of quantizing and dequantizing activation tensors using GranQ.
In Algorithm 1, both the scaling and quantization steps are implemented using vectorized tensor operations across all channels. Specifically, the per-channel minimum and maximum values are computed simultaneously, and the resulting scaling factors and zero-points are applied through broadcasting. This eliminates the need for channel-wise loops and allows efficient execution on modern hardware accelerators.
In contrast, Algorithm 2 follows a conventional channel-wise quantization approach, where the scaling parameters are computed sequentially for each channel using an explicit for-loop. The quantization and dequantization steps are still vectorized per channel. However, the scaling computation remains scalar and iterative. This leads to increased computational overhead, especially when applied to deep networks with many channels. By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
Algorithm 1 GranQ: Vectorized Per-Channel Scaling for Activation Quantization
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
@@ A.2 Algorithm
4:    Reshape:
$A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
$\triangleright$ Flatten spatial dimensions
5:    Decompose by channel:
$\vec{A}_{\min}=\min_{h,w}(A_{l}),\quad\vec{A}_{\max}=\max_{h,w}(A_{l})$
$\triangleright$ Extract per-channel range
6:    Compute scaling parameters:
$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$
7:    Quantize (vectorized):
$A_{q,l}=\left\lfloor\frac{A_{l}}{\vec{s}}+\vec{z}\right\rceil$
8:    Dequantize (vectorized):
$A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$
@@
10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
Algorithm 2 Conventional Channel-Wise Activation Quantization (Scalar-loop for scaling)
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:    Compute per-channel scaling parameters:
5:    for each channel $c$ in $C$ do
$s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1},\quad z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
6:    end for
7:    Apply quantization (vectorized):
$A_{q,l}[c,:,:]=\left\lfloor\frac{A_{l}[c,:,:]}{s_{c}}+z_{c}\right\rceil,\quad\forall c$
8:    Dequantize (vectorized):
$A_{\text{deq},l}[c,:,:]=s_{c}(A_{q,l}[c,:,:]-z_{c}),\quad\forall c$
9:end for
10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
A.3 Details of Experimental Setup
A.3.1 Datasets
This paper outlines the experimental settings for CIFAR-10 [13], CIFAR-100 [13], and ImageNet (ILSVRC-2012) [14] datasets. CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images. ImageNet comprises 1,000 classes, with 1.2 M training and 50 K validation images, all resized to 224 × 224 for our experiments. We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
A.3.2 Experimental Setup
The data generation process followed the settings of AdaDFQ [1]. Specifically, ACGAN [2] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [1]. For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD [47] as the optimizer, with momentum 0.9 and weight decay 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch [3], and visualizations were generated with matplotlib [4]. All experiments were conducted on an NVIDIA RTX 3090 GPU.
A.3.3 Comparison Methods
  • GDFQ [5]: GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation. It demonstrated that effective quantization is possible even without access to original data.
  • AIT [6]: AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation. It also provided a systematic analysis of training strategies in the ZSQ domain.
  • AdaDFQ [1]: AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
  • TexQ [7]: TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
  • RIS [8]: RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
  • GenQ [10]: GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
  • AKT [11]: AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
  • SynQ [12]: SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
These methods cover various aspects of ZSQ, including data generation, training objectives. We compare them to demonstrate the effectiveness and generality of GranQ, which focuses on improving the quantization process itself by introducing an optimized, vectorized scaling mechanism. This structural enhancement allows GranQ to be seamlessly integrated into existing QAT-based ZSQ pipelines, providing consistent gains across different settings. Furthermore, as several methods have not yet released official code, we expect that GranQ can be readily applied as their implementations become publicly available.
A.4 Extra Analysis
A.4.1 Data Distribution Analysis
@@
Figure 6: Pixel distribution comparison between (a) original and (b) synthetic inputs on ImageNet. The synthetic inputs, generated by AdaDFQ [17], exhibit higher skewness and a rightward shift, indicating the presence of more extreme activation values and motivating the need for granular scaling in quantization.
Figure 6 compares the pixel distributions of original and synthetic inputs derived from ImageNet. The skewness of the original input is 0.2599, whereas that of the synthetic input increases to 0.4877, with the distribution shifted toward higher pixel values. This indicates that synthetic data, often generated without real image priors, tends to contain a greater number of extreme activation values. As a result, the likelihood of outliers increases, which fixed-range quantization cannot effectively handle. In particular, when a single global range is applied across all channels, the presence of outliers can compress the majority of values into a narrow dynamic range, leading to severe quantization distortion. These observations underscore the limitations of conventional scaling strategies. To address this, we propose a granular, channel-wise scaling approach that dynamically adapts to per-channel statistics and better preserves activation fidelity under distributional mismatch.
A.4.2 Computational Efficiency Analysis
Although GranQ and conventional channel-wise quantization share a higher theoretical complexity of $\mathcal{O}(L\cdot C\cdot H\cdot W)$ compared to the layer-wise baseline with $\mathcal{O}(L\cdot H\cdot W)$, all three methods apply their respective scaling strategies only during the training phase. Since QAT fixes the scaling factors as static parameters prior to deployment, these operations are not executed during inference, resulting in no runtime overhead. Consequently, the per-channel granularity used in GranQ does not incur any additional inference latency. As shown in Table 7, all methods achieve comparable inference latency regardless of the scaling strategy, demonstrating that GranQ maintains runtime efficiency while benefiting from fine-grained scaling during training.
Table 7: Comparison of activation quantization latency on ResNet-18 (ImageNet, batch size = 128) under a 3-bit quantization setting. All latency values are measured per batch and averaged over 3 evaluations, with standard deviation reported in parentheses. Theoretical complexity reflects quantization-time operations and does not affect inference latency, which is measured separately.
Method Theoretical Complexity Scaling Strategy Inference Latency (ms)
Activation (Layer-wise) $O(L\cdot H\cdot W)$ Layer-wise uniform scaling 44.194 (+0.190)
Activation (Channel-wise, Scalar-loop) $O(L\cdot C\cdot H\cdot W)$ Channel-wise scalar scaling (loop-based) 44.272 (+0.158)
GranQ (Channel-wise, Vectorization) $O(L\cdot C\cdot H\cdot W)$ Channel-wise vectorized scaling 44.231 (+0.186)
A.4.3 Analysis of Activation Preservation

### v5

@@ Introduction
Table 5: Ablation study on the impact of pre-scaling and vectorization in quantization. ✓indicates whether vectorization is applied. Latency is measured with a batch size of 16 using AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
@@
While fine-grained activation quantization often incurs high computational cost due to per-channel scaling, GranQ maintains efficiency by vectorizing the scaling computation. By reshaping activation tensors, our method eliminates the overhead of iterative channel-wise operations. To assess this in practice, we measured quantization latency across various batch sizes (16, 32, 64, 128, and 200). Figure 4 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels.
@@
The two channel-wise methods perform per-channel quantization. The conventional method computes scaling factors sequentially, whereas GranQ pre-applies them in a vectorized manner across channels, enabling pre-scaling for efficient parallel accumulation. A key observation from Figure 4 is that GranQ achieves substantial accuracy improvement with only a minimal latency overhead compared to conventional layer-wise quantization. Furthermore, Table 2 and Figure 4 shows that GranQ substantially reduces quantization error and more effectively preserves activation information. These results demonstrate that GranQ enables fine-grained quantization with significantly improved efficiency.
Additionally, GranQ demonstrates that per-channel quantization can be efficiently executed by vectorizing scaling factor computation. As shown in Table 5, it achieves a quantization latency of 0.236 ms, comparable to the layer-wise baseline (0.227 ms), and significantly outperforms the conventional channel-wise method (103.671 ms) suffering from scalar-loop overhead.
Discussion
Why GranQ is Effective?
@@ Discussion
In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. These limitations become more severe in low-bit settings, especially when synthetic data exhibit distributional shifts. To overcome these issues, GranQ pre-computes per-channel scaling in a vectorized form, enabling efficient parallel accumulation. By integrating scaling into the quantization phase, we removed runtime scaling during accumulation, allowing precise and efficient parallel execution.
Limitation and Future-work
In ZSQ, synthetic data often exhibit broader and noisier activation distributions (see Figure 6 in appendix). While GranQ effectively handles this in QAT through dynamic per-channel scaling, its static form is less suited for PTQ. Future work will explore adapting GranQ to better support PTQ.
Conclusion
Layer-wise and channel-wise quantization are both commonly used in weight quantization, as weights remain fixed after training. In contrast, it is harder to apply channel-wise activation quantization because input-dependent activations make fine-grained scaling difficult. These issues are especially critical in ZSQ, with data synthesized from FP statistics. In such scenarios, synthetic activations often deviate from real data, amplifying quantization error and complicating accurate scaling. In addition, channel-wise quantization poses a computational burden due to the need for per-channel scaling during accumulation. Since scaling factors differ across channels, accumulation requires separate multiplications, limiting parallelism and increasing latency. This overhead discourages the use of channel-wise methods in practice, despite their higher representational fidelity. To address these issues, we propose GranQ, which vectorizes scaling to enable fine-grained quantization and efficient accumulation across channels. Through extensive experiments, we demonstrate that GranQ achieves notable accuracy and efficiency. Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
@@ Appendix A Appendix
Notation
We provide an overview of the main symbols and definitions used throughout the paper, along with their precise mathematical formulations.
@@
AA Original activation tensor
$\vec{A}_{q}$ Quantized activation vector (per channel)
$\vec{\tilde{A}}$ Pre-scaled activation
$\vec{A}_{\min},\vec{A}_{\max}$ Per-channel min and max activation values
$\vec{s}$ Per-channel scaling factor vector
$\vec{z}$ Per-channel zero-point vector
@@
$C,H,W$ Channel, Height, Width dimensions of activation
$\mathbb{R}^{C\times H\times W}$ Activation tensor space
$\lfloor\cdot\rceil$ Rounding function
$\cdot$ Element-wise multiplication
$\sum_{c=1}^{C}$ Channel-wise summation
$w_{c}$ Weight for channel $c$
$y_{l}$ Output after accumulation at layer $l$
@@
Scaling Factor $\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1}$
Zero-point $\vec{z}=\left\lfloor\frac{-\vec{A}_{\min}}{\vec{s}}\right\rceil$
Forward Quantization $\vec{A}_{q}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil$
Pre-scaled Dequantization $\vec{\tilde{A}}=\vec{s}\cdot(\vec{A}_{q}-\vec{z})$
Output Accumulation $y=\sum_{c=1}^{C}w_{c}\cdot\vec{\tilde{A}}_{c}$
@@
Table 6: Notation and expressions for vectorized channel-wise activation quantization with pre-scaling.
Details of Experimental Setup
Datasets
This paper outlines the experimental settings for CIFAR-10 (Krizhevsky, Hinton et al. 2009), CIFAR-100 (Krizhevsky, Hinton et al. 2009), and ImageNet (ILSVRC-2012) (Deng et al. 2009) datasets. CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images. ImageNet comprises 1,000 classes, with 1.2 M training and 50 K validation images, all resized to 224 × 224 for our experiments. We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
Experimental Setup
The data generation process followed the settings of AdaDFQ (Qian et al. 2023). Specifically, ACGAN (Odena, Olah, and Shlens 2017) was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ (Qian et al. 2023). For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD (Nesterov 1983) as the optimizer, with momentum 0.9 and weight decay 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch (Paszke et al. 2019), and visualizations were generated with matplotlib (Hunter 2007). All experiments were conducted on an NVIDIA RTX 3090 GPU.
Comparison Methods
  • GDFQ (Xu et al. 2020): GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation. It demonstrated that effective quantization is possible even without access to original data.
  • AIT (Choi et al. 2022): AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation. It also provided a systematic analysis of training strategies in the ZSQ domain.
  • AdaDFQ (Qian et al. 2023): AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
  • TexQ (Chen et al. 2023): TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
  • RIS (Bai et al. 2024): RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
  • GenQ (Li et al. 2024): GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
  • AKT (Hong et al. 2024): AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
  • SynQ (Kim, Kim, and Kang 2025): SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
Algorithm 2 Conventional Channel-Wise Quantization (Scalar-loop for scaling)
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$, weights $\{W_{l}\}_{l=1}^{L}$
2:Output vectors $\{y_{l}\}_{l=1}^{L}$
@@
4:  $A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
5:  for each channel $c$ in $C$ do
6:   $s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1}$
7:   $z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
8:   $A_{q,l}[c,:]=\left\lfloor\frac{A_{l}[c,:]}{s_{c}}+z_{c}\right\rceil$
@@
10:  Compute output:
$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\left(s_{c}\cdot(A_{q,l}[c,:]-z_{c})\right)$
11:end for
12:return $\{y_{l}\}_{l=1}^{L}$
Algorithm
While prior works apply vectorized computation only to the quantization step, they compute per-channel scaling parameters through iterative operations. GranQ also vectorizes this scaling step, enabling fully parallel quantization. Algorithm 1 summarizes the process with pre-scaling.
In Algorithm 1, the entire quantization procedure is implemented using vectorized tensor operations across all channels, enabling efficient accumulation without runtime scaling. Specifically, the per-channel minimum and maximum values are computed in parallel, and the corresponding scaling factors and zero-points are applied via broadcasting. This removes the need for channel-wise loops and enables efficient accumulation without runtime scaling, making the procedure highly suitable for modern hardware accelerators.
In contrast, Algorithm 2 adopts a conventional channel-wise quantization scheme, where scaling parameters are computed iteratively for each channel using a for-loop. Although the quantization and dequantization steps are applied in a vectorized manner per channel, the overall process remains sequential due to scalar-based scaling. This incurs significant computational overhead, particularly in networks with a large number of channels.
By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
Extra Analysis
Data Distribution Analysis
@@
Figure 5: Pixel distribution comparison between (a) original and (b) synthetic inputs on ImageNet. Synthetic inputs generated by AdaDFQ (Qian et al. 2023) exhibit a rightward shift with higher skewness, indicating more extreme activations and highlighting the need for granular scaling.
Figure 5 compares the pixel distributions of original and synthetic inputs derived from ImageNet. The skewness of the original input is 0.2599, whereas that of the synthetic input increases to 0.4877, with the distribution shifted toward higher pixel values. This indicates that synthetic data, often generated without real image priors, tends to contain a greater number of extreme activation values. As a result, the likelihood of outliers increases, which fixed-range quantization cannot effectively handle. In particular, when a single global range is applied across all channels, the presence of outliers can compress the majority of values into a narrow dynamic range, leading to severe quantization distortion. These observations underscore the limitations of conventional scaling strategies.
Analysis of Activation Preservation